# Discovery of Seven Cold and Distant Brown Dwarfs with JWST RUBIES

### v1

@@ 5.2 Spectral Variations Among Distant T Dwarfs
The mid- and late-type T dwarfs that have been identified in deep JWST/NIRSpec spectral surveys extend out to 1–2 kpc, a distance that encompasses a diversity of populations and corresponding ages and compositions. As such, these early discoveries provide the opportunity to examine the influence of secondary parameters such as surface gravity and metallicity on cool brown dwarf spectra. To illustrate this, Figure 7 compares the NIRSpec prism spectra of five distant mid- and late-type T dwarfs, including RUBIES-BD-1 (T7), -6 (T8), and -7 (T5), the T6 dwarf UNCOVER-BD-2 (estimated distance of 2.1–2.6 kpc; Burgasser et al. 2024), and the T6 dwarf JADES-GS-BD-9 (estimated distance of 1.8–2.3 kpc; Hainline et al. 2024b). All of these sources are similarly normalized to focus on spectral morphology, and are compared to the JWST/NIRSpec prism spectrum of the T6 dwarf standard SDSS J1624+0029 (Beiler et al., 2024a). We separately evaluate the 0.9–2.4 m NIR and 3.6–5.0 m MIR regions to focus on key spectral features. In the NIR, we observe subtle variations between these spectra in several key features, including the depth of the 1.15 m and 1.6 m CH4 bands (shallowest in RUBIES-BD-7, deepest in RUBIES-BD-1), the width of the 1.25 m peak (widest in RUBIES-BD-7, narrowest in JADES-GS-BD-9 and RUBIES-BD-6), and the shape and brightness of the 2.1 m peak (broad in RUBIES-BD-7, shifted redward in UNCOVER-BD-2, shifted blueward in RUBIES-BD-1, flat in JADES-GS-BD-9 and RUBIES-BD-6). In the MIR, the broad 4 m peak shows a variety of shapes among these sources, from a relatively featureless and blueshifted peak in the spectra of RUBIES-BD-7 and UNCOVER-BD-2, to the more structured peaks in the spectra of JADES-GS-BD-9 and RUBIES-BD-7, the latter showing strong CO2 and CO absorption bands at 4.3 m and 4.6 m, also observed in the spectrum of SDSS J1624+0029.
@@
To contextualize these variations, we compare in the right panels of Figure 7 Elf Owl atmosphere models with variations in temperature, surface gravity, metallicity, and vertical mixing efficiency relative to a baseline model appropriate for a normal thin disk T6 dwarf (Teff = 800 K, = 4.5, [M/H] = 0, and = 4). NIR variations sensitive to these features have been discussed extensively in the literature (e.g., Burgasser et al. 2002; Leggett et al. 2010; Pinfield et al. 2012; Zhang et al. 2017a; Mukherjee et al. 2024; Burgasser et al. 2025b). Here, we highlight that variations in the depths of CH4 bands, the width of the 1.25 m peak, and shape of the 2.1 m peak are all driven primarily by temperature and metallicity variations in the models. Surface gravity plays a greater role in setting the relative brightness of the 1.05 m and 1.6 m peaks, while vertical mixing plays only a minor role in shaping the 2.1 m peak. In the MIR, the models show that metallicity plays a greatest role in shaping the 4 m peak, with low metallicity models missing the CO2 and CO bands and producing a relatively featureless broad plateau. Vertical mixing also modulates the CO2 feature. Temperature primarily shifts the location of the 4 m peak (blueward for warmer temperatures), while surface gravity only modestly influences the depth of the CO band.
@@
This analysis suggests that all of these parameters can drive the variations observed among these similarly-classified distant T dwarfs. As such, a comparison of multiple features is necessary to isolate metallicity variations from other secondary parameters. There is an important caveat to this analysis, in that current models do not fully capture the complex atmosphere chemistry among the observed molecular constituents, Most importantly, the 4.3 m CO2 band is known to be far stronger than can be explained even with efficient vertical mixing (Yamamura et al., 2010; Beiler et al., 2024b). We also note that none of these spectra show clear evidence of PH3 absorption that was initially claimed in the distant (0.6–1.2 kpc) late T dwarf UNCOVER-BD-3 (Burgasser et al., 2024, see Beiler et al. 2024b for further discussion) and more recently detected in the metal-poor T dwarf companion Wolf 1130C (Burgasser et al., 2025a). The absence or weakness of this feature in most brown dwarf and exoplanet spectra has been a challenge to chemical models (Morley et al., 2018; Rowland et al., 2024).
@@ 6 Summary
We compared flux-calibrated spectra to three sets of atmosphere models: Sonora Elf Owl, Sonora Diamondback, and SAND. We find the L dwarf spectra are best matched to the Diamondback and SAND models, likely due to the presence of condensates in their atmospheres; the early-T dwarf spectra are best matched to the SAND models; and the mid- and late-T dwarf spectra are best matched to the Elf Owl and Diamondback models, the latter with efficient sedimentation or no clouds. Spectra of the two L dwarfs, RUBIES-BD-2 and -3, show the most significant deviations from the models, possibly due to incorrect treatment of condensates or the failure to incorporate reddening corrections, as previously noted by Tu et al. (2025a). Only one source, RUBIES-BD-5, shows significant evidence of being metal-poor based on the model fits.
@@
Examination of the present sample of mid- and late-type T dwarfs found in deep JWST spectral surveys at distances of 1–2 kpc show subtle but distinct variations in the strengths of major molecular bands and the shapes of inter-band peaks, most notably the 2.1 m flux peak and the 4.3 m CO2 band. Examination of comparable Elf Owl models point to multiple factors that may be responsible for these variations, requiring further examination to isolate the expected metallicity variations in these distant objects.
This study adds to the growing sample of deep brown dwarf spectra from JWST/NIRSpec prism data collected in deep survey programs such as CEERS (Finkelstein et al., 2023), UNCOVER (Bezanson et al., 2024), JADES (D’Eugenio et al., 2025), CANUCS (Willott et al., 2022), BoRG (Roberts-Borsani et al., 2025), NEXUS (Shen et al., 2024), and CAPERS (Dickinson et al., 2024), among others (cf. Tu et al. 2025b). These discoveries increase the diversity of physical parameters sampled among known brown dwarfs, and provide the opportunity to improve atmosphere and evolutionary models of low-temperature atmospheres more broadly, not only for thick disk and halo brown dwarf but also candidate brown dwarfs identified in globular clusters (Gerasimov et al., 2024) and stellar stream populations (Burgasser et al., 2025b), as well as gas giant exoplanets orbiting metal-poor stars (Barbato et al., 2019; Alves et al., 2022). Building larger and more diverse samples of cool brown dwarfs from these surveys can also facilitate their use to address broader questions in Milky Way science, such as the evolution and distribution of chemical abundances, the Milky Way’s star formation and accretion history, and the dynamical history of various Galactic populations (Burgasser, 2004a; Ryan et al., 2017; Aganze et al., 2022b; Best et al., 2024; Raghu et al., 2024; Honaker & Gizis, 2025).
@@ Appendix A Near-Infrared Spectral Standards
Table 7 lists the near-infrared spectral standards used for the classification of our sources in Section 2.3. Dwarf standards are defined in Burgasser et al. (2006) and Kirkpatrick et al. (2010); subdwarf standards are defined in Zhang et al. (2017b, 2018a); Greco et al. (2019); and Burgasser et al. (2025b). We drew from previously-published data acquired with several different instruments, which were resampled to the wavelength-dependent resolution of the JWST/NIRSpec prism mode (Birkmann et al., 2022). The resampled spectra are included as part of the online materials for the article.
@@
References. — [1] Burgasser & McElwain (2006); [2] Bardalez Gagliuffi et al. (2014); [3] Burgasser et al. (2007); [4] Burgasser (2007); [5] Chiu et al. (2006); [6] Reid et al. (2006); [7] Burgasser et al. (2006); [8] Looper et al. (2007); [9] Burgasser & Splat Development Team (2017); [10] Burgasser et al. (2004); [11] Burgasser et al. (2006); [12] Kirkpatrick et al. (2010); [13] Burgasser et al. (2025b); [14] Zhang et al. (2018b); [15] Greco et al. (2019); [16] Zhang et al. (2019); [17] Pinfield et al. (2014); [18] Zhang et al. (2017b); [19] Schmidt et al. (2010); [20] Meisner et al. (2021); [21] Kirkpatrick et al. (2021); [22] Burgasser et al. (2010); [23] Burgasser (2004b); [24] Burgasser et al. (2009); [25] Schneider et al. (2020); [26] Lodieu et al. (2022).

### v2

@@ 5.2 Spectral Variations Among Distant T Dwarfs
The mid- and late-type T dwarfs that have been identified in deep JWST/NIRSpec spectral surveys extend out to 1–2 kpc, a distance that encompasses a diversity of populations, ages, and compositions. As such, these early discoveries provide the opportunity to examine the influence of secondary parameters such as surface gravity and metallicity on cool brown dwarf spectra. To illustrate this, Figure 7 compares the NIRSpec/Prism spectra of five distant mid- and late-type T dwarfs, including RUBIES-BD-1 (T7), -6 (T8), and -7 (T5), the T6 dwarf UNCOVER-BD-2 (estimated distance of 2.1–2.6 kpc; 2024ApJ...962..177B), and the T6 dwarf JADES-GS-BD-9 (estimated distance of 1.8–2.3 kpc; 2024ApJ...975...31H). All of these sources are similarly normalized to focus on spectral morphology, and are compared to the NIRSpec/Prism spectrum of the T6 dwarf standard SDSS J1624+0029 (2024ApJ...973..107B). We separately evaluate the 0.9–2.4 m NIR and 3.6–5.0 m MIR regions to focus on key spectral features. In the NIR, we observe subtle variations between these spectra in several key features, including the depth of the 1.15 m and 1.6 m CH4 bands (shallowest in RUBIES-BD-7, deepest in RUBIES-BD-1), the width of the 1.25 m peak (widest in RUBIES-BD-7, narrowest in JADES-GS-BD-9 and RUBIES-BD-6), and the shape and brightness of the 2.1 m peak (broad in RUBIES-BD-7, shifted redward in UNCOVER-BD-2, shifted blueward in RUBIES-BD-1, flat in JADES-GS-BD-9 and RUBIES-BD-6). In the MIR, the broad 4 m peak shows a variety of shapes among these sources, from a relatively featureless and blueshifted peak in the spectra of RUBIES-BD-7 and UNCOVER-BD-2, to the more structured peaks in the spectra of JADES-GS-BD-9 and RUBIES-BD-7, the latter showing strong CO2 and CO absorption bands at 4.3 m and 4.6 m, also observed in the spectrum of SDSS J1624+0029.
@@
To contextualize these variations, we compare in the right panels of Figure 7 Elf Owl atmosphere models with variations in temperature, surface gravity, metallicity, and vertical mixing efficiency relative to a baseline model appropriate for a thin disk T6 dwarf (Teff = 800 K, = 4.5, [M/H] = 0, and = 4). NIR variations sensitive to these features have been discussed extensively in the literature (e.g., 2002ApJ...564..421B; 2010ApJ...710.1627L; 2012MNRAS.422.1922P; 2017MNRAS.468..261Z; 2024ApJ...963...73M; 2025ApJ...982...79B). Here, we highlight that variations in the depths of CH4 bands, the width of the 1.25 m peak, and shape of the 2.1 m peak are all driven primarily by temperature and metallicity variations in the models. Surface gravity plays a greater role in setting the relative brightness of the 1.05 m and 1.6 m peaks, while vertical mixing plays only a minor role in shaping the 2.1 m peak. In the MIR, the models show that metallicity plays greatest role in shaping the 4 m peak, with low metallicity models missing the CO2 and CO bands and producing a relatively featureless broad plateau. Vertical mixing also modulates the CO2 feature. Temperature primarily shifts the location of the 4 m peak (blueward for warmer temperatures), while surface gravity only modestly influences the depth of the CO band.
@@
This analysis suggests that all of these parameters can drive the variations observed among these similarly-classified distant T dwarfs. As such, a comparison of multiple features is necessary to isolate metallicity variations from other secondary parameters. There is an important caveat to this analysis, in that current models do not fully capture the complex atmosphere chemistry among the observed molecular constituents, Most importantly, the 4.3 m CO2 band is known to be far stronger than can be explained even with efficient vertical mixing (2010ApJ...722..682Y; 2024ApJ...973...60B). We also note that none of these spectra show clear evidence of PH3 absorption that was initially claimed in the distant (0.6–1.2 kpc) late T dwarf UNCOVER-BD-3 (2024ApJ...962..177B, see 2024ApJ...973...60B for further discussion) and more recently detected in the metal-poor T dwarf companion Wolf 1130C (burgasser-ph3). The absence or weakness of this feature in most brown dwarf and exoplanet spectra a challenge to chemical models (2018ApJ...858...97M; 2024ApJ...977L..49R).
@@ 6 Summary
We compared flux-calibrated spectra to three sets of atmosphere models: Sonora Elf Owl, Sonora Diamondback, and SAND. We find the L dwarf spectra are best matched to the Diamondback and SAND models, likely due to the presence of condensates in their atmospheres; the early-T dwarf spectra are best matched to the SAND models; and the mid- and late-T dwarf spectra are best matched to the Elf Owl and Diamondback models, the latter with efficient sedimentation or no clouds. Spectra of the two L dwarfs, RUBIES-BD-2 and -3, show the most significant deviations from the models, possibly due to incorrect treatment of condensates or the failure to incorporate reddening corrections, as previously noted by 2025ApJ...980..230T. Only one source, RUBIES-BD-5, shows significant evidence of being metal-poor based on the model fits.
@@
Examination of the present sample of mid- and late-type T dwarfs found in deep JWST spectral surveys at distances of 1–2 kpc show subtle but distinct variations in the strengths of major molecular bands and the shapes of inter-band peaks, most notably the 2.1 m flux peak and the 4.3 m CO2 band. Examination of comparable Elf Owl models point to multiple factors that may be responsible for these variations, requiring further examination to isolate the these distant objects.
This study adds to the growing sample of deep brown dwarf spectra from JWST/NIRSpec Prism data collected in deep survey programs such as CEERS (2023ApJ...946L..13F), UNCOVER (2024ApJ...974...92B), JADES (2025ApJS..277....4D), CANUCS (2022PASP..134b5002W), BoRG (2025ApJ...983...18R), NEXUS (2024arXiv240812713S), and CAPERS (2024jwst.prop.6368D), among others (cf. 2025arXiv251002026T). these discoveries increase the diversity of physical parameters sampled among known brown dwarfs, and provide the opportunity to improve atmosphere and evolutionary models of low-temperature atmospheres more broadly, not only for thick disk and halo brown dwarf but also candidate brown dwarfs identified in globular clusters (2024ApJ...971...65G) and stellar stream populations (2025ApJ...982...79B), as well as gas giant exoplanets orbiting metal-poor stars (2019A&A...621A.110B; 2022MNRAS.517.4447A). Building larger and more diverse samples of cool brown dwarfs from these surveys can also facilitate their use to address broader questions in Milky Way science, such as the evolution and distribution of chemical abundances, the Milky Way’s star formation and accretion history, and the dynamical history of various Galactic populations (2004ApJS..155..191B; 2017ApJ...847...53R; 2022ApJ...934...73A; 2024ApJ...967..115B; 2024ApJ...974..222R; 2025ApJ...985...48H).
@@ Appendix A Near-Infrared Spectral Standards
Table 7 lists the near-infrared spectral standards used for the classification of our sources in Section 2.3. Dwarf standards are defined in Burgasser_2006 and 2010ApJS..190..100K; subdwarf standards are defined in 2017MNRAS.464.3040Z; 2018MNRAS.479.1383Z; 2019AJ....158..182G; and 2025ApJ...982...79B. We drew from previously-published data acquired with several different instruments, which were resampled to the wavelength-dependent resolution of the JWST/NIRSpec prism mode (Birkmann_2022). The resampled spectra are included as part of the online materials for the article.
@@
References. — [1] 2006AJ....131.1007B; [2] 2014ApJ...794..143B; [3] 2007ApJ...658..557B; [4] 2007ApJ...659..655B; [5] 2006AJ....131.2722C; [6] 2006AJ....132..891R; [7] 2006ApJ...637.1067B; [8] 2007AJ....134.1162L; [9] 2017ASInC..14....7B; [10] 2004AJ....127.2856B; [11] 2006ApJ...639.1095B; [12] 2010ApJS..190..100K; [13] 2025ApJ...982...79B; [14] 2018MNRAS.480.5447Z; [15] 2019AJ....158..182G; [16] 2019MNRAS.486.1260Z; [17] 2014MNRAS.437.1009P; [18] 2017MNRAS.464.3040Z; [19] 2010AJ....139.1045S; [20] 2021ApJ...915..120M; [21] 2021ApJS..253....7K; [22] 2010AJ....139.2448B; [23] 2004ApJ...614L..73B; [24] 2009ApJ...697..148B; [25] 2020ApJ...898...77S; [26] 2022AA...663A..84L.